suppose that you invest $6000 for your money at an annual intrest rate of of 4% and invest the rest of your money in another accont earning annual intrest rate of 11%. if the two investment earn a total of $1065 for the year , find the total amaount of money that you have invested.
Please help me solve

Answer by checkley77(12844)

6,000*.04+.11x=1,065
240+.11x=1,065
.11x=1,065-240
.11x=825
x=825/.11
x=7,500 ans. for the amount invested @ 11%.
6,000+7,500=$13,500 is the total amount invested.
Proof:
6,000*.04+7,500*.11=1,065
240+825=1,065
1,065=1,065